A particle of charge 3 x 〖10〗^(-9) C is located at point (0, 0) in a certain coordinate system. Assuming all lengths are in meters, calculate the potential at the following points A (3, 3) and B (0, 2). How much work is needed to take a particle of charge 2 x 〖10〗^(-9) C from point A to point B?
1
Expert's answer
2020-05-29T09:48:02-0400

Potential created by a point charge "q" at distance "r" from it is "\varphi = k \frac{q}{r}", where "k = \frac{1}{4 \pi \epsilon_0} = 9 \cdot 10^9 N \cdot m^2 \cdot C^{-2}" is Coulomb's constant.

Point A is at distance "r_A = \sqrt{3^2 + 3^2} = 3 \sqrt{2}" from charge at "(0,0)", and point B is at distance "r_B = 2".

Hence, "\varphi_A = \frac{9 \cdot 10^9 N m^2 C^{-2} \cdot 3 \cdot 10^{-9} C}{3 \sqrt{2} m} \approx 6.4 V" and "\varphi_B = \frac{9 \cdot 10^9 N m^2 C^{-2} \cdot 3 \cdot 10^{-9} C}{2 m} \approx 13.5 V".

Work needed to take charge "q = 2 \cdot 10^{-9} C" from point A to point B is "A = q (\varphi_A - \varphi_B) \approx -1.42 \cdot 10^{-8} J".